OpenStudy (anonymous):

Use newton's method to approximate the solution to x^5-2=3x+1; find the third iteration using x1=1 as the initial guess.

OpenStudy (anonymous):

\[x_2=x_1-\frac{f\left(x_1\right)}{f'\left(x_1\right)}\]

OpenStudy (anonymous):

Rewrite function as

OpenStudy (anonymous):

\[x^5-3x-3=0\]

OpenStudy (anonymous):

\[1-\frac{\left(x^5-3x-3\right)}{5x^4-3}\]

OpenStudy (anonymous):

Continueplz

OpenStudy (anonymous):

Lets' plug in 1 \[1-\frac{\left(1^5-3(1)-3\right)}{5(1)^4-3}\] =7/2

OpenStudy (anonymous):

Now we have to do whole thing again with 7/2 instead of 1

OpenStudy (anonymous):

So is 2.8 my final ans?

OpenStudy (anonymous):

I got 2.28